xx WaveToy1DF77

WaveToy1DF77

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11amplitude10Scope: privatex01  REAL

11Description: Amplitude of initial Gaussian wave

11 Range 10  x01  Default: 1.0

110:*  10

11bound 10Scope: privatex01  KEYWORD

11Description: Which boundary condition to apply

11 Range 10 x01 Default: zero
11zero  10Set boundary values to zero
11radiation 10Radiation/wave boundary conditions
11none  10Don’t apply any condition

11

11center 10Scope: privatex01  REAL

11Description: x-position of center of Gaussian wave

11 Range 10 x01 Default: 0.5
11*:*  10Anywhere

11

11sigma 10Scope: privatex01  REAL

11Description: Width of Gaussian wave

11 Range 10 x01 Default: 0.02
11*:*  10Anything

11

7 Interfaces

General Implements: wavetoy1d

Inherits: boundary

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  evolve1d  10  x01  compact110
11                 10  phi  x01 description11The evolved scalar field
11                 10  x01 dimensions111
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL
11  initial  10  x01  compact110
11                 10  timederiv  x01 description11For initial data
11                 10  x01 dimensions111
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  dummy  10  x01  compact110
11                 10  phixtemp  x01 description11Just for creating initial data
11                 10  x01 dimensions111
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  coords1d  10  x01  compact110
11                 10  x  x01 description11The 1D x-coordinate
11                 10  x01 dimensions111
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/WaveToy1DF77. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11evolve1d[3] coords1d11x

11x  11x

11

Scheduled Functions

CCTKxSTARTUP x wavetoy1df77xstartup x register wavetoy1df77 banner x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxWRAGH x wavetoy1df77xregistercoords x register wavetoy1df77 coordinates x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxBASEGRID x wavetoy1df77xcoord x set up 1d coordinates x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxINITIAL x wavetoy1df77xinitial x initial data for 1d wave equation x

11x11Language:11fortran
11 x 11Storage: 11initial
11 x 11x  11dummy
11 x 11Type: 11function

11

CCTKxEVOL x wavetoy1df77xevolve x evolution of 1d wave equation x

11x11Language:11fortran
11 x 11Sync:  11evolve1d
11 x 11Type: 11function

11

CCTKxEVOL x wavetoy1df77xboundary x apply boundary conditions x

11x11After: 11wavetoy1df77xevolve
11 x 11Language:11fortran
11 x 11Type: 11function

Aliased Functions x

11Alias Name: xxxxxxx11Function Name:
11WaveToy1DF77xCoord 11SpatialCoordinates